A grocer mixes walnuts selling at 55 cents a pound and almonds selling at 70 cents a pound, making a mixture of 30 pounds to sell at 61 cents. How many pounds of walnuts should he use? Represent the walnuts with x and the almonds with y. transale the word statement in 2 equations. I tried this for about 30 minutes. so far i got x + y = 30 and 55x + 70y = 610.
Answer by stanbon(75887)

Quantity: x + y = 30
Value: 55x + 70y = 61*30
-----
Multiply thru Quantity by 55 to get:
55x + 55y = 55*30
55x + 70y = 61*30
-----
Subtract and solve for "y":
15y = 6*30
y = 12 lbs. (amt of 55 cent walnuts needed)
---
Solve for "x":
x + y = 30
y = 18 lbs. (amt of 70 cent almonds needed)
==============
Cheers,
Stan H.
=========================